# Radiative corrections to the $\rm S,T,U$ parameters and their impact on the $W$ boson mass in the 331 model

### v1

@@ 1 Introduction
In addition to the discrepancy in the prediction of the $W$ boson mass reported by the CDF, there are various theoretical and experimental motivations to explore BSM. Numerous models have been proposed to address these concerns. One such model is the 331 extension of the SM, which is based on the concept of higher weak isospin symmetry, denoted by $SU(3)_{L}$, directly extending from $SU(2)_{L}$ [5, 6, 7, 8, 9]. Notably, this novel gauge framework offers potential solutions to key questions concerning fermion family number [5], the quantization of electric charge [10], discrepancies in the third quark family [11], and strong ${\cal CP}$ conservation [12]. Moreover, it has the potential to address issues such as dark matter [13, 14], flavor physics [15], neutrino mass generation [16], cosmic inflation, and the baryon asymmetry [17, 18, 19].
Various 331 extensions have been proposed in the literature. Among them, the minimal 331 model [20] and the 331 model with right-hand neutrinos represent two typical examples [5, 6, 7]. Additionally, several other intriguing models have emerged, including the economical 331 model[21], the reduced 331 model [20], the simple 331 model [22], the flipped 331 model[23], the 331 model with exotic charged leptons [24], and the 331 model with neutral (heavy) fermions [25].
@@ 2.1 Scalar Sector
[…]
@@ 3.1 The ${\rm S,T,U}$ parameters
Here, $\hat{\alpha}(M_{Z})$ denotes the renormalized electromagnetic fine-structure constant evaluated at the $Z$-boson mass scale, whereas $s_{W}\equiv\sin\theta_{W}$ and $c_{W}\equiv\cos\theta_{W}$ represent the sine and cosine of the weak mixing angle, respectively. The vacuum polarization functions $\Pi^{\text{new}}_{VV^{\prime}}(q^{2})$ encode the contributions from BSM to the gauge boson two-point functions, with the subscripts $VV^{\prime}$ referring to specific gauge boson combinations: $\Pi^{\text{new}}_{WW}$ for the $W$ bosons, $\Pi^{\text{new}}_{ZZ}$ for the $Z$ bosons, $\Pi^{\text{new}}_{Z\gamma}$ for the $Z$–photon mixing, and $\Pi^{\text{new}}_{\gamma\gamma}$ for the photon self-energy. The latest values for these parameters[2] are given by:
@@ 4.2 Input parameters
The parameter $f$ is expressed as $f\equiv kv_{3}$, where typically, the constant $k$ is assumed to be around unity. However, the masses of the Higgs bosons exhibit sensitivity to the magnitude of $f$. This sensitivity leads to large values for the ${\rm S,T,U}$ parameters, consequently imposing constraints on the value of $f$, as will be explained in the following section. Additionally, the mass of the light Higgs boson, $M_{h}$, is highly sensitive to the input parameters, particularly the couplings $\lambda_{1}$ and $\lambda_{2}$. Consequently, we only consider points that result in $M_{h}$ falling within the range of 123-127 GeV to allow for theoretical uncertainties from unknown higher-order corrections and model-dependent effects.
@@ 4.3 Higgs Boson Masses
In this section, we present our results concerning the masses of Higgs bosons plotted on the $f-v_{3}$ plane in Figs. 2-7. Since $M_{h}$ is independent of $f$ and $v_{3}$, we do not present the predictions for $M_{h}$. Fig. 2 displays the value of $M_{H_{2}}$ in $f-v_{3}$ plane. The left plot showcases our findings without imposing ${\rm S,T,U}$ constraints, while the right plot displays results where only points satisfying ${\rm S,T,U}$ experimental constraints within the $3\sigma$ range are included. The color bar indicates the magnitude of $M_{H_{2}}$. As expected, the value of $M_{H_{2}}$ is influenced by both parameters, $f$ and $v_{3}$. It can be observed that the ${\rm S,T,U}$ constraints preclude values of $v_{3}$ below $1500\,\,\mathrm{GeV}$ and above $2300\,\,\mathrm{GeV}$. In the left plot, the dominance of green and yellow points in the region where $v_{3}>2000\,\,\mathrm{GeV}$ and $f<3\,\,\mathrm{GeV}$ indicates $M_{H_{2}}>1000\,\,\mathrm{GeV}$. However, the right plot reveals that this region is excluded by the ${\rm S,T,U}$ parameters. As anticipated based on Eq. (8), $M_{H_{W}}$ and $M_{A}$ exhibit nearly degenerate behavior. Consequently, a comparable trend to that of $M_{H_{2}}$ is observed for $M_{H_{W}}$ and $M_{A}$, as can be seen in Fig. 3 and Fig. 4 respectively.
@@
Figure 4: The 331 predictions for $M_{A}$ with (right plot) and without (left plot) ${\rm S,T,U}$ constraints in the $f-v_{3}$ plane. The color bar represents the values of $M_{h}$.
@@
Figure 6: The 331 predictions for $M_{H_{V}}$ with (right plot) and without (left plot) ${\rm S,T,U}$ constraints in the $f-v_{3}$ plane. The color bar represents the values of $M_{H_{Y}}$.
@@ 4.4 Gauge Boson Masses
In this section, we present our results for gauge boson masses, specifically the $M_{Z^{\prime}}$, $M_{Y^{\pm}}$, and $M_{V^{\pm\pm}}$. Although we do not anticipate significant contributions from the $Y^{\pm}$ and $V^{\pm\pm}$ gauge bosons to the ${\rm S,T,U}$ parameters due to their nearly degenerate masses, it is still useful to determine their possible mass ranges considering the influence of the ${\rm S,T,U}$ parameters. In Fig. 8, we present 331 predictions for $M_{Z^{\prime}}$ on the $f-v_{3}$ plane, illustrated in both the right and left plots with and without ${\rm S,T,U}$ constraints, respectively. The color bar indicates the values of $M_{Z^{\prime}}$. It is evident that the ${\rm S,T,U}$ constraints restrict the value of $M_{Z^{\prime}}$ to nearly $7\,\,\mathrm{TeV}$, contrasting with its potential range up to $13\,\,\mathrm{TeV}$ in their absence. Additionally, the values of $M_{Y^{\pm}}$ and $M_{V^{\pm\pm}}$ are similarly bounded to approximately $750\,\,\mathrm{GeV}$ by the ${\rm S,T,U}$ constraints, as evidenced by Fig. 9 and Fig. 10.
@@ 5 Conclusions
M. E. G. research is supported by the Spanish MICINN, under grant PID2022-140440NB-C22.

### v2

@@ Abstract
Radiative corrections to the $\rm S,T,U$ parameters and their
impact on the $W$ boson mass in the 331 model
M. Rehman1***email: m.rehman@comsats.edu.pk, M.A. Iqbal1email: mradeeliqbal@yahoo.com, M.E. Gómez2email: mario.gomez@dfa.uhu.esand  O. Panella3§§§email: orlando.panella@cern.ch
1Department of Physics, Comsats University Islamabad, 44000 Islamabad, Pakistan
2Dpt. de Ciencias Integradas y Centro de Estudios Avanzados en Física Matemáticas y Computación, Campus del Carmen, Universidad de Huelva, Huelva 21071, Spain
3 INFN, Sezione di Perugia, Via A. Pascoli, I-06123, Perugia, Italy
@@ 1 Introduction
In addition to the discrepancy in the prediction of the $W$ boson mass reported by the CDF, there are various theoretical and experimental motivations to explore BSM. Numerous models have been proposed to address these concerns. One such model is the 331 extension of the SM, which is based on the concept of higher weak isospin symmetry, denoted by $SU(3)_{L}$, directly extending from $SU(2)_{L}$ [5, 6, 7, 8, 9]. Notably, this novel gauge framework offers potential solutions to key questions concerning fermion family number [5], the quantization of electric charge [10, 11], discrepancies in the third quark family [12], and strong $\cal CP$} conservation [13, 14]. Moreover, it has the potential to address issues such as dark matter [15, 16, 17], flavor physics [18, 19, 20], neutrino mass generation [21, 22, 23], cosmic inflation, and the baryon asymmetry [24, 25, 26].
Various 331 extensions have been proposed in the literature. Among them, the minimal 331 model [27] and the 331 model with right-hand neutrinos represent two typical examples [5, 6, 7]. Additionally, several other intriguing models have emerged, including the economical 331 model[28], the reduced 331 model [27], the simple 331 model [29], the flipped 331 model[30], the 331 model with exotic charged leptons [31], and the 331 model with neutral (heavy) fermions [32]. While certain realizations of the 331 model are known to develop a Landau pole at relatively low energies [33], the  Refs. [34, 35] discuss possible ways to avoid this problem.
@@ 2.1 Scalar Sector
The Higgs boson $h$ is identifiable as the particle detected at the Large Hadron Collider (LHC)[52, 53] at CERN, with a mass of $125\,\text{GeV}$. Supposedly, $H_{2}$ and $H_{3}$ are heavier. $h$ and $H_{2}$ emerge from $\rho$ and $\eta$, while $H_{3}$ arises from $\chi$, as a consequence of the two-step symmetry breaking. Additionally, within the Higgs sector of the 331 model, there are three charged scalars denoted as $H_{W}^{\pm}$, $H_{V}^{\pm\pm}$, and $H_{Y}^{\pm}$. The tree-level masses of the Higgs bosons within the 331 model can be summarized as:
@@ 3.1 The $\rm S,T,U$} parameters
Here, $\hat{\alpha}(M_{Z})$ denotes the renormalized electromagnetic fine-structure constant evaluated at the $Z$-boson mass scale, whereas $s_{W}\equiv\sin\theta_{W}$ and $c_{W}\equiv\cos\theta_{W}$ represent the sine and cosine of the weak mixing angle, respectively. The vacuum polarization functions $\Pi^{\text{new}}_{VV^{\prime}}(q^{2})$ encode the contributions from BSM to the gauge boson two-point functions, with the subscripts $VV^{\prime}$ referring to specific gauge boson combinations: $\Pi^{\text{new}}_{WW}$ for the $W$ bosons, $\Pi^{\text{new}}_{ZZ}$ for the $Z$ bosons, $\Pi^{\text{new}}_{Z\gamma}$ for the ZZ–photon mixing, and $\Pi^{\text{new}}_{\gamma\gamma}$ for the photon self-energy. The latest values for these parameters[2] are given by:
@@ 4.2 Input parameters
We restrict $v_{3}$ to values up to $5000\,\,\mathrm{GeV}$ to avoid the Landau pole[43]. The parameter $f$ is expressed as $f\equiv kv_{3}$, where typically, the constant $k$ is assumed to be around unity. However, the masses of the Higgs bosons exhibit sensitivity to the magnitude of $f$. This sensitivity leads to large values for the $\rm S,T,U$} parameters, consequently imposing constraints on the value of $f$, as will be explained in the following section. Additionally, the mass of the light Higgs boson, $M_{h}$, is highly sensitive to the input parameters, particularly the couplings $\lambda_{1}$ and $\lambda_{2}$. Consequently, we only consider points that result in $M_{h}$ falling within the range of 123-127 GeV to allow for theoretical uncertainties from unknown higher-order corrections and model-dependent effects.
@@ 4.3 Higgs Boson Masses
In this section, we present our results concerning the masses of Higgs bosons plotted on the $f-v_{3}$ plane in Figs. 2-7. Since $M_{h}$ is independent of $f$ and $v_{3}$, we do not present the predictions for $M_{h}$. Fig. 2 displays the value of $M_{H_{2}}$ in $f-v_{3}$ plane. The left plot showcases our findings without imposing $\rm S,T,U$} constraints, while the right plot displays results where only points satisfying $\rm S,T,U$} experimental constraints within the $3\sigma$ range are included. The color bar indicates the magnitude of $M_{H_{2}}$. As expected, the value of $M_{H_{2}}$ is influenced by both parameters, $f$ and $v_{3}$. It can be observed that the $\rm S,T,U$} constraints preclude values of $v_{3}$ below $1500\,\,\mathrm{GeV}$ and above $2300\,\,\mathrm{GeV}$. In the left plot, the dominance of green and yellow points in the region where $v_{3}>2000\,\,\mathrm{GeV}$ and $f>3\,\,\mathrm{GeV}$ indicates $M_{H_{2}}>1000\,\,\mathrm{GeV}$. However, the right plot reveals that this region is excluded by the $\rm S,T,U$} parameters. As anticipated based on Eq. (8), $M_{H_{W}}$ and $M_{A}$ exhibit nearly degenerate behavior. Consequently, a comparable trend to that of $M_{H_{2}}$ is observed for $M_{H_{W}}$ and $M_{A}$, as can be seen in Fig. 3 and Fig. 4 respectively.
@@
Figure 4: The 331 predictions for $M_{A}$ with (right plot) and without (left plot) $\rm S,T,U$} constraints in the $f-v_{3}$ plane. The color bar represents the values of $M_{A}$.
@@
Figure 6: The 331 predictions for $M_{H_{Y}}$ with (right plot) and without (left plot) $\rm S,T,U$} constraints in the $f-v_{3}$ plane. The color bar represents the values of $M_{H_{Y}}$.
@@ 4.4 Gauge Boson Masses
In this section, we present our results for gauge boson masses, specifically the $M_{Z^{\prime}}$, $M_{Y^{\pm}}$, and $M_{V^{\pm\pm}}$. Although we do not anticipate significant contributions from the $Y^{\pm}$ and $V^{\pm\pm}$ gauge bosons to the $\rm S,T,U$} parameters due to their nearly degenerate masses, it is still useful to determine their possible mass ranges considering the influence of the $\rm S,T,U$} parameters. In Fig. 8, we present 331 predictions for $M_{Z^{\prime}}$ on the $f-v_{3}$ plane, illustrated in both the right and left plots with and without $\rm S,T,U$} constraints, respectively. The color bar indicates the values of $M_{Z^{\prime}}$. It is evident that the $\rm S,T,U$} constraints restrict the value of $M_{Z^{\prime}}$ to nearly $7\,\,\mathrm{TeV}$, contrasting with its potential range up to $13\,\,\mathrm{TeV}$ in their absence.
The analysis in Refs. [58, 59] place lower bounds on $M_{Z^{\prime}}$ that can be translated into constraints on the 331 symmetry-breaking scale $v_{3}$ via the relation given in Eq. (15), valid in the decoupling limit $v_{3}\gg v_{1,2}$. The experimental mass limit $M_{Z^{\prime}}\gtrsim 4\ \mathrm{TeV}$ quoted there corresponds to the value at which the theoretical prediction for $\sigma(pp\to Z^{\prime})\times\mathrm{BR}(Z^{\prime}\to\ell\ell)$ meets the experimental upper bound. Among other factors, this prediction depends sensitively on the $Z^{\prime}$ couplings to quarks and leptons. These couplings control both the production rate, which at LHC energies is dominated by light-quark parton luminosities and scales approximately as $\sigma(pp\to Z^{\prime})\propto(g_{Z^{\prime}}^{q})^{2}$, and the leptonic branching ratio,
$\mathrm{BR}(Z^{\prime}\to\ell\ell)=\frac{\Gamma(Z^{\prime}\to\ell\ell)}{\Gamma_{\rm tot}},\qquad\Gamma(Z^{\prime}\to f\bar{f})\propto(g_{Z^{\prime}}^{f})^{2}M_{Z^{\prime}}.$
Here, $g_{Z^{\prime}}^{f}$ denotes the effective coupling of the $Z^{\prime}$ to a generic fermion $f$. Any non-uniform change between quark and lepton couplings modifies both the production rate and the leptonic branching fraction, while the opening or closing of additional decay channels alters $\Gamma_{\rm tot}$ and hence $\mathrm{BR}(Z^{\prime}\to\ell\ell)$. Since the fermion charge assignments—and therefore the coupling factors $g_{Z^{\prime}}^{f}$—differ among 331 variants, in particular between $\beta=\pm 1/\sqrt{3}$ and $\beta=-\sqrt{3}$, a faithful translation of the published $\beta=\pm 1/\sqrt{3}$ limit into our $\beta=-\sqrt{3}$ setup requires recomputing $\sigma\times\mathrm{BR}$ with the appropriate charge assignments, while adopting the same acceptance and K-factor assumptions as in the experimental analyses. In practice, the $\beta=\pm 1/\sqrt{3}$ bound is expected to be somewhat stronger. In the absence of a dedicated recast, and allowing for plausible $\mathcal{O}(1)$ variations in the couplings and width effects, we conservatively assign a model-dependent uncertainty of 12TeV1\!-\!2\ $\mathrm${TeV}, under which the limit remains compatible with our results.
Additionally, the values of $M_{Y^{\pm}}$ and $M_{V^{\pm\pm}}$ are similarly bounded to approximately $750\,\,\mathrm{GeV}$ by the $\rm S,T,U$} constraints, as evidenced by Fig. 9 and Fig. 10. We note that a recent collider recast analysis[60] of ATLAS multi-lepton data sets a lower limit of about $M_{V^{\pm\pm}}\gtrsim 1.3\,\,\mathrm{TeV}$ (90% C.L.) under the assumption that the doubly charged vector boson decays almost exclusively into same-sign SM lepton pairs. At face value, this lower bound appears in tension with the upper bound $M_{V^{\pm\pm}}\lesssim 750\,\,\mathrm{GeV}$ inferred from our analysis. However, several important caveats should be emphasized. First, the collider bound depends sensitively on the branching ratio into SM leptons: if new decay channels of $V^{\pm\pm}$ into exotic fermions or scalar states are kinematically open, the effective leptonic branching fraction is reduced and the ATLAS limit correspondingly weakens. Second, the STU constraints are themselves highly sensitive to the scalar-sector quartic couplings. While the requirement $M_{h}\simeq 125$ GeV tightly constrains combinations of $\lambda_{1}$ and $\lambda_{2}$, other quartic couplings (notably $\lambda_{3}$ and the primed $\lambda^{\prime}_{i3}$) remain free within our scan and directly control the heavy-scalar masses and their isospin splittings. Modest variations of these couplings can both reduce the $T$ parameter by making heavy scalar multiplets more degenerate and open additional cascade decay channels for $V^{\pm\pm}$. These effects can substantially relax the apparent tension between the STU-derived upper limit and collider-derived lower bound. A dedicated combined study of oblique parameters, scalar potential parameters, and collider signatures is therefore required before any definitive claims.
@@ 5 Conclusions
M. E. G. research is supported by the Spanish MICINN, under grant PID2022-140440NB-C22. O. P. is supported by the INFN grant ENP (Exploring New Physics).